two planes leave an airport at 12:00 noon . If the east-bound plane flies at 560 mph and the west-bound plane flies at 500 mph, at what time will they be 2,000 miles apart ?
so far i have. 560+500=2000
1060=2000
i don't know where to go from there

Answer by checkley77(12844)

D=RT
2,000=(560+500)T
2,000=1,060T
T=2,000/1,060
T=1.887 HOURS. OR 1 HOUR & 53 MINUTES.
12:00+1:53=1:53 PM. THEY WILL BE 2,000 MILES APART.
PROOF:
2,000=(560+500)1.887
2,000=1060*1.887
2,000=2,000